OpenStudy (anonymous):

limit of (x-1)^2(x+1) as x approaches 3

OpenStudy (zzr0ck3r):

is there any reason you just cant plug in 3?

OpenStudy (alexwee123):

^^ ya substitute 3 for x

OpenStudy (anonymous):

i dont think. there's any steps i would have to show?

OpenStudy (anonymous):



OpenStudy (anonymous):

do you mean this?

OpenStudy (zzr0ck3r):

note: unless ding so makes you devide by zero, take a sqrt( a negative number ), or take a ln( negative number ), then you can plug in for x

OpenStudy (zzr0ck3r):

you can show what ever step you would in algebra.

OpenStudy (zzr0ck3r):

(3-1)^2 * (3+1) = 2^2 * 4 .....

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2}(x-1)^2(x+1)\]

OpenStudy (anonymous):

sorry not 2 but 3

OpenStudy (zzr0ck3r):

3 or 2? Not that it makes a dif here, just plug in for x

OpenStudy (zzr0ck3r):

ahh ...

OpenStudy (zzr0ck3r):

now if it where lim x->3 of 1/(3-x) then you could not just plug in for x because you would get 1/0

OpenStudy (zzr0ck3r):

were

OpenStudy (anonymous):

find the indicated one-sided limit. if the limiting value is inifinte indicate whther is is \[+\infty or -\], \[\lim_{x \rightarrow 4} (3x-9)\]